# Aligned Novel View Image and Geometry Synthesis via Cross-modal Attention Instillation

### v2

@@ 3.2 Aligned novel view geometry generation
We perform novel view geometry prediction alongside image synthesis using an identical architecture to the image generation architecture: a geometry denoising network (U-Net) that predicts a target view pointmap, paired with its geometry reference network that receives reference view pointmaps $\{P_{n}\}_{n=1}^{N}$ as input. Similarly to Ke et al. [13], our geometry denoising network and geometry reference network are fine-tuned from an image denoising U-Net to predict pointmap modalities instead of images, detailed in Sec. A of the appendix. As described in lower half of Fig. 2, the geometry generation architecture predicts a pointmap for the target viewpoint $\pi_{t}$, identical to the image denoising network, and therefore receives the same target and reference correspondence conditions, $\mathbf{c}^{\mathbf{t}}$ and $\mathbf{c}^{\mathbf{r}}$, as the image generation architecture. This encourages the pointmap generated by the geometry generation network to align with the generated image.
Figure 3: Per-modality attention visualization.
However, we find this naïve approach of giving identical correspondence condition to be insufficient in achieving alignment between the generated image and its geometry. Instead, the two modalities exhibit different behaviors when completing same regions, with geometry denoising significantly outperforming image inpainting, as demonstrated in Fig. 3. This disparity can be attributed to geometry completion being inherently more deterministic due to the coarse nature of the modality, making it substantially easier to learn than image inpainting. As shown in Fig. 3(a-b), when completing a partially visible wheel, the image denoising model, shown in (a), fails to establish such correspondences, the geometry prediction model (b) correctly attends to other wheel locations on the object for structural reference, generating higher-accuracy completion result.
In this manner, in extrapolative scenarios where large portions of the target view are unobserved, the image generation pipeline encounters significant challenges due to the severe ambiguity from partial observations. On the other hand, while geometry completion is more deterministic due to the coarse nature of its representations, this coarseness also restricts the available information for accurate completion, limiting its accuracy and performance overall.
@@
Cross-modal attention instillation. In this light, we propose cross-modal attention instillation, where the spatial attention maps for the geometry prediction U-Net is substituted by the attention maps from image denoising U-Net to achieve synergistic effects. Specifically, the key and query features extracted from the image denoising U-Net are leveraged by the spatial attention layers of the geometry denoising U-Net. Let $K^{I}$ and $Q^{I}$ denote the image key and query features from the image U-Net, respectively, and let $V^{P}=[V^{P}_{t},V^{P}_{1},V^{P}_{2},\dots V^{P}_{N}]$ represent the geometry value features from the geometry U-Net that generates pointmap $P_{t}$. The spatial attention is as follows:
@@
where $d_{k}$ is the dimensionality of the key features.
This method offers several key advantages. As shown in Fig. 3, injecting the attention map across the modalities reinforces the alignment between generated images and their geometries. The image denoising U-Net receives deterministic training signals from the geometry completion network, which regularizes its generation process, yielding enhanced consistency and inpainting capability, as demonstrated in Fig. 3 (c). The geometry prediction U-Net also leverages rich semantic features from the image domain to achieve more accurate geometry completion. Additionally, as the attention maps solely serve as structural information for aggregating value features, our architecture prevents the harmful mixing of features from different modalities that is often observed in cross-modal architectures [7].
@@ 3.3 Proximity-based mesh conditioning
The off-the-shelf geometry models [29, 13] typically generate a sparse 3D point cloud with noise and errors, which becomes particularly severe when the target viewpoint deviates significantly from the reference viewpoints. Erroneous projections from the sparse point cloud causes misalignment in the generation process, as the networks cannot differentiate valids from erroneous signals, this compromises the consistency and accuracy of generated images and their geometry.
To address these challenges, we propose proximity-based mesh conditioning. We begin by converting the sparse point cloud into a mesh representation using ball-pivoting algorithm [1], which reduces erroneous projections and yields denser, interpolated conditions for the generation networks. Therefore, instead of employing the naïve projected pointmap $P^{\Pi}_{t}$ as the correspondence condition, we utilize the pointmap derived from the projected mesh, denoted as $X^{\Pi}_{t}$, as our correspondence condition.
Furthermore, we augment the correspondence condition with the mesh’s depth and normal map, enabling the network to prioritize reliable correspondences while filtering out noise and erroneous projections. Specifically, we channel-wisely concatenate the partial depth map $D^{\Pi}_{t}$ and normal map $N^{\Pi}_{t}$ acquired from our converted mesh to the correspondence condition embedding. Accordingly, our final correspondence conditions are expressed as follows:
$\mathbf{c}^{\mathbf{t}}=[\mathcal{E}(X^{\Pi}_{t}),D^{\Pi}_{t},N^{\Pi}_{t},M_{t}],$
$\mathbf{c}^{\mathbf{r}}_{n}=[\mathcal{E}(X_{n}),D_{n},N_{n},\mathbf{1}].$ (8)
@@ 4 Experiment
4.1 Implementation details
For image denoising networks, we initialize from Stable Diffusion 2.1 [24]. Reference networks share identical architecture but omit timestep embeddings, serving only for feature extraction. We train on RealEstate10K [34], Co3D [22], and MVImgNet [33] using pseudo ground-truth geometry from VGGT [28]. During training, only reference pointmaps are used for warping and proximity-based mesh conditioning. At inference, VGGT predicts camera poses and pointmaps for projection to target viewpoints. For geometry denoising networks, we initialize from Marigold [13]—specifically the normal prediction head due to its three-channel similarity to pointmaps. We repurpose Marigold to generate complete geometry from incomplete warped RGB conditioning rather than target images.
@@ 4.1 Implementation details
We define extrapolative camera settings as cases when the target camera position ${\bm{t}}_{t}$ lies outside the convex hull of reference camera positions of $\{{\bm{t}}_{n}\}_{n=1}^{N}$, ignoring viewing direction for simplicity. Alternatively, if ${\bm{t}}_{t}$ cannot be expressed as a convex combination $\Sigma_{n=1}^{N}\alpha_{n}{\bm{t}}_{n}$ where $\alpha_{n}\geq 0$ and $\Sigma_{n=1}^{N}\alpha_{n}=1$. This geometric constraint means that significant portions of the target view may contain regions that are either occluded in reference views or lie beyond the observable scene boundaries. Unlike interpolative settings where novel views can be reconstructed through geometric triangulation and feature matching between overlapping reference regions, extrapolative synthesis requires generating plausible scene content based on learned priors about scene structure and appearance, therefore requiring generative priors of diffusion models.
Figure 4: Qualitative results. We demonstrate our qualitative results on the Co3D [22] dataset, conducting feed-forward novel view synthesis while generating aligned geometry in a robust and consistent manner.
@@ 4.2 Experimental results
Results on Co3D. Figure 4 presents a qualitative demonstration of our approach on the Co3D dataset, where our model reconstructs a novel view image with its aligned geometry from three reference images (i.e., 3-view setting). The generated images are consistent with the references, and their geometry is accurately aligned with the synthesized views. Visualizations of the pointmap from multiple viewpoints reveal that the generated geometry aligns well without requiring additional scale-and-shift fitting, due to our formulation of the depth prediction task as a depth completion problem combined with image inpainting. Note that the novel view images and their geometries are directly generated from our denoising network at target viewpoints, requiring no additional optimization of NeRF or 3DGS to acquire geometry or synthesize unconstrained target view synthesis, overcoming the limitations of previous diffusion-based novel view synthesis models [6, 30, 26].
@@
2-view PixelSplat [3] 14.66 0.517 0.334 12.75 0.329 0.637
MVSplat [4] 12.22 0.416 0.423 13.94 0.473 0.385
NoPoSplat [31] 13.58 0.393 0.545 14.04 0.414 0.530
@@
1-view LucidDreamer [5] 11.14 0.423 0.440 12.09 0.481 0.419
GenWarp [25] 9.85 0.315 0.527 9.54 0.298 0.538
@@
Table 1: Zero-shot quantitative comparison. We compare our model to existing feedforward NVS methods (2-view setting) and warping-and-inpainting methods (1-view setting) on DTU [34] dataset, which is zero-shot setting for all the models. Our method showing superior performance in both extrapolative and original (interpolative) setting in both single-view and stereo-view settings.
Figure 5: Qualitative comparison with inpainting method on DTU [34] dataset. Our qualitative comparison with the naive warping-and-inpainting method demonstrates our model’s zero-shot generalization capabilities to unseen data, as well as its ability to robustly handle erroneous warped geometries for geometrically consistent generation.
Results on DTU. In Table 1 and Fig.5, we compare against feedforward methods (PixelSplat[3], MVSplat [4], DUSt3R [29], NopoSplat [31]) using two reference views, and warping-and-inpainting methods using single reference images (LucidDreamer [5], GenWarp [25]). We evaluate on DTU dataset [11], which was not used in training in all the models, demonstrating zero-shot generalization. For fair comparison, all warping-and-inpainting methods use the same geometry prediction model [28], VGGT. Since our model excels in extrapolative settings, we introduce an extrapolative view selection that samples the furthest target cameras in the dataset. Our method achieves state-of-the-art performance in both extrapolative and interpolative settings regardless of viewpoint count. Qualitative results in Fig. 5 show our model effectively filters artifacts and noise from predicted point clouds during warping, producing clean, complete images with well-aligned geometry. In contrast, naive inpainting of warped images yields numerous artifacts and inconsistent features, demonstrating our model’s robust handling of erroneous warped conditions.
@@
PixelSplat [3] 14.01 0.582 0.384 23.85 0.806 0.185
MVSplat [4] 12.13 0.534 0.380 23.98 0.811 0.176
NoPoSplat [31] 14.36 0.538 0.389 25.03 0.838 0.160
@@
Table 2: In-domain comparison. We compare our model to existing feedforward NVS methods in Realestate10K [34] dataset, our method showing superior performance in extrapolative setting.
@@
Results on RealEstate10K. In Table 2, we compare our method against feedforward few-shot novel view synthesis approaches—PixelSplat [3], MVSplat [4], DUSt3R [29], and NopoSplat [31]—on RealEstate10K [34]. Similarly, we evaluate both interpolative (target near reference views) and extrapolative (target in unobserved locations) conditions using two reference views. For extrapolation, we sample reference images from the latter third of video frames and targets from the first third, exploiting the forward camera progression in source videos. Our model outperforms feedforward methods in extrapolative settings while maintaining competitive interpolative performance. As shown in Figure 6, conventional feedforward models struggle with large missing areas due to limited generative capabilities, whereas our approach effectively infers missing geometry and generates plausible imagery for unseen regions while preserving observed region fidelity. Notably, we can realistically inpaint partially visible objects (e.g., sofas, railings) with well-aligned geometry synthesis. We attribute this to geometric awareness from attention distillation, which regularizes structural attention maps and enhances generation.
Figure 7: Comparison with LVSM.
Comparison against LVSM [12]. We compare our method against LVSM [12], a leading feedforward approach for novel view synthesis. As shown in Fig. 7, our method demonstrates superior performance in extrapolative scenarios where significant portions of the target view are unobserved in the reference images. While LVSM accurately reconstructs regions with sufficient geometric overlap from reference views, it fails to extrapolate beyond the observable scene boundaries, producing blurry and inconsistent content in occluded areas. In contrast, our generative approach leverages learned scene priors to synthesize plausible content for these missing regions, maintaining sharp detail and geometric consistency throughout the entire target view. This fundamental difference highlights the limitation of purely reconstructive methods in extrapolative settings, where the lack of reference information necessitates generative capabilities rather than geometric interpolation.
@@ 4.3 Ablation
(b) (a) + Pointmap condition 16.93 0.594 0.243
(c) (b) + Proximity-based mesh 17.01 0.601 0.238
(d) (c) + Cross-modal instillation 17.41 0.614 0.229
@@
Component analysis.
We conduct quantitative ablation studies on our method using the RealEstate10K dataset under the extrapolative setting. Our results in Table 3 demonstrate the contribution of each component. The baseline (a) receives no geometric conditioning, while (b) introduces naive pointmap conditioning. When we add proximity-based mesh conditioning (c) and cross-modal attention distillation (d), we observe progressive performance improvements.
Figure 8: Analysis on number of reference viewpoints. Our model’s multi-view aggregated attention enables our model to generalize towarding receiving arbitrary number of reference images.
@@
Analysis on the number of input viewpoints. Our model conducts aggregated attention to generate novel views from reference images, it can receive an arbitrary number of input viewpoints for generation, as an additional reference image corresponds to simply concatenating additional reference viewpoint’s features within our attention architecture. To demonstrate this, in Figure 8 and Table 5, we increase the number of reference viewpoints for a model trained at 2-viewpoint setting and analyze its effects in both image quality and geometric accuracy: the results demonstrate even without being trained on the given number of inputs, our model benefits strongly from additional viewpoints, showing the generalization capability of our aggregated attention architecture to various number of input reference viewpoints.

### v3

@@ 3.2 Aligned novel view geometry generation
We perform novel view geometry prediction alongside image synthesis using the same architecture as in image generation. The geometry denoising network (U-Net) predicts a target view pointmap, paired with the geometry reference network that receives reference view pointmaps $\{P_{n}\}_{n=1}^{N}$. Similar to Ke et al. (2024), our geometry denoising and reference networks are fine-tuned from image denoising U-Nets to predict pointmaps instead of images (ref., Appx.˜A in the Appendix). As described in the lower half of Fig.˜2, the geometry generation part predicts a pointmap for the target viewpoint $\pi_{t}$, as in the image denoising network, receiving the same conditions, $\mathbf{c}^{\mathbf{t}}$ and $\mathbf{c}^{\mathbf{r}}$. This makes the predicted pointmaps generated by the geometry generation network align with the generated image.
Figure 3: Effects of cross-modal attention instillation.
However, we found this naïve approach of giving identical conditions was insufficient for the alignment between the generated image and its geometry. The two modalities exhibit different behaviors when completing void regions, with geometry denoising outperforming image inpainting, as in Fig. 3. This performance difference can be attributed to the more deterministic nature of geometry completion than image generation, as geometric tasks have stronger structural constraints and less inherent ambiguity, resulting in more consistent and robust outcomes.
As shown in Fig. 3(a–b), when completing a partially visible wheel, the image denoising (a) fails to establish such correspondences, whereas the geometry prediction (b) correctly attends to other wheel locations for structural reference, producing more realistic completion. Conversely, while the geometry denoising network is better at geometry completion, the spatial attention within these networks struggles to get fine-grained cross-viewpoint correspondences due to the lack of semantic cues, as in (b), where the geometry network’s attention is diffusely distributed in comparison to the more focused attention in (a). This complementary relationship motivates a synergistic framework where the geometry network leverages semantic cues from image features via their spatial attention map to achieve more detailed geometry generation, while the structural constraints from geometry completion implicitly guide the image denoising network toward more robust and consistent inpainting.
@@
Cross-modal attention instillation. In this light, we propose cross-modal attention instillation, where the spatial attention maps for the geometry prediction U-Net are substituted by the attention maps from the image denoising U-Net to achieve synergistic effects. Specifically, the key and query features extracted from the image denoising U-Net are leveraged by the spatial attention layers of the geometry denoising U-Net. Let $K^{I}$ and $Q^{I}$ denote the image key and query features from the image U-Net, respectively, and let $V^{P}=[V^{P}_{t},V^{P}_{1},V^{P}_{2},\dots V^{P}_{N}]$ represent the geometry value features from the geometry U-Net that generates pointmap $P_{t}$. The spatial attention is as follows, with $d_{k}$ as the dimensionality of the key features:
@@
This method offers several key advantages. As shown in Fig. 3, injecting the attention map across the modalities reinforces the alignment between generated images and their geometries. The image denoising U-Net receives deterministic training signals from the geometry completion network, which regularizes its generation process, yielding enhanced consistency and inpainting capability, as demonstrated in Fig. 3 (c). The geometry prediction U-Net also leverages rich semantic features from the image domain to achieve more accurate geometry completion. Additionally, since attention maps serve only as structural cues for aggregating value features, our architecture avoids the detrimental cross-modal feature mixing often observed in the prior work (He et al., 2024).
@@ 3.3 Proximity-based mesh conditioning
The off-the-shelf geometry models (Wang et al., 2024; Ke et al., 2024) typically generate a sparse 3D point cloud with noise and errors, which becomes particularly severe when the target viewpoint deviates significantly from the reference viewpoints. Erroneous projections from the sparse point cloud cause misalignment in the generation process, as the networks cannot differentiate valid from erroneous, and harm the accuracy of generated images and their geometry.
To address these challenges, we propose proximity-based mesh conditioning. We convert the sparse point cloud into a mesh representation with the ball-pivoting algorithm (Bernardini et al., 1999), which reduces erroneous projections and yields dense projections for the generation networks. Therefore, instead of employing the naïve projected pointmap $P^{\Pi}_{t}$ as the correspondence condition, we utilize the pointmap derived from the projected mesh, denoted $X^{\Pi}_{t}$, as our correspondence condition.
Furthermore, we augment the correspondence condition with the mesh’s depth and normal map, enabling the network to prioritize reliable correspondences while filtering out noise and erroneous projections. Specifically, we channel-wise concatenate the partial depth map $D^{\Pi}_{t}$ and normal map $N^{\Pi}_{t}$ acquired from our converted mesh to the correspondence condition embedding. Accordingly, our final correspondence conditions are expressed as follows:
$\mathbf{c}^{\mathbf{t}}=[\mathcal{E}(X^{\Pi}_{t}),D^{\Pi}_{t},N^{\Pi}_{t},M_{t}],\quad\mathbf{c}^{\mathbf{r}}_{n}=[\mathcal{E}(X_{n}),D_{n},N_{n},\mathbf{1}].$ (8)
@@ 4 Experiment
4.1 Implementation and experimental details
For image denoising networks, we initialize from Stable Diffusion 2.1 (Rombach et al., 2022). Reference networks share identical architecture but omit timestep embeddings, serving only for feature extraction. We train on RealEstate10K (Zhou et al., 2018), Co3D (Reizenstein et al., 2021), and MVImgNet (Yu et al., 2023) using pseudo ground-truth geometry from VGGT (Wang et al., 2025). During training, only reference pointmaps are used for warping and proximity-based mesh conditioning. At inference, VGGT predicts reference camera poses and pointmaps for projection to target viewpoints. For geometry denoising networks, we initialize from Marigold (Ke et al., 2024);s normal prediction, finetuning it to generate complete geometry from incomplete warped RGB conditioning rather than target images.
@@ 4.1 Implementation and experimental details
We define extrapolative camera settings as cases when the target camera position ${\bm{t}}_{t}$ lies outside the convex hull of reference camera positions of $\{{\bm{t}}_{n}\}_{n=1}^{N}$, ignoring viewing direction for simplicity. Alternatively, if ${\bm{t}}_{t}$ cannot be expressed as a convex combination $\Sigma_{n=1}^{N}\alpha_{n}{\bm{t}}_{n}$ where $\alpha_{n}\geq 0$ and $\Sigma_{n=1}^{N}\alpha_{n}=1$. This geometric constraint means that significant portions of the target view may contain regions that are either occluded in reference views or lie beyond the observable scene boundaries, which requires the model to generate plausible scene content based on learned priors about scene structure and appearance to fill in the unknown regions.
Figure 4: Qualitative results. We demonstrate our qualitative results on the Co3D (Reizenstein et al., 2021) dataset, conducting NVS while generating aligned geometry robustly and consistently.
@@ 4.2 Experimental results
Results on Co3D. Figure 4 demonstrates our approach on Co3D, reconstructing novel views with aligned geometry from three reference images. Generated images maintain consistency with references while achieving accurate geometric alignment. Multi-viewpoint pointmap visualizations show well-aligned geometry without scale-and-shift fitting, enabled by formulating depth prediction as completion combined with inpainting. Our denoising network directly generates novel views and geometry at target viewpoints without additional NeRF or 3DGS optimization, overcoming limitations of prior diffusion-based methods (Gao et al., 2024; Wu et al., 2023; Shi et al., 2023).
@@
2-view PixelSplat (Charatan et al., 2024) 14.66 0.517 0.334 12.75 0.329 0.637
MVSplat (Chen et al., 2024) 12.22 0.416 0.423 13.94 0.473 0.385
NoPoSplat (Ye et al., 2024) 13.58 0.393 0.545 14.04 0.414 0.530
@@
1-view LucidDreamer (Chung et al., 2023) 11.14 0.423 0.440 12.09 0.481 0.419
GenWarp (Seo et al., 2024) 9.85 0.315 0.527 9.54 0.298 0.538
@@
Table 1: Zero-shot quantitative comparison. We compare our model to existing feedforward NVS methods (2-view setting) and warping-and-inpainting methods (1-view setting) on DTU Zhou et al. (2018) dataset, which is zero-shot setting for all the models. Our method shows superior performance in both extrapolative and original (interpolative) setting in both single-view and stereo-view settings.
Figure 5: Qualitative comparison with inpainting method on DTU (Zhou et al., 2018) dataset. Our qualitative comparison with the naive warping-and-inpainting method demonstrates our model’s zero-shot generalization capabilities to unseen data, as well as its ability to robustly handle erroneous warped geometries for geometrically consistent generation.
Zero-shot results on DTU. Table 1 and Fig. 5 compare our method against feedforward approaches (PixelSplat (Charatan et al., 2024), MVSplat (Chen et al., 2024), DUSt3R (Wang et al., 2024), NopoSplat (Ye et al., 2024)) using two views, and warping-inpainting methods using single views (LucidDreamer (Chung et al., 2023), GenWarp (Seo et al., 2024)). Evaluation on DTU (Jensen et al., 2014) demonstrates zero-shot generalization capability of our method. For fair comparison, warping methods use identical geometry prediction (VGGT (Wang et al., 2025)). We introduce extrapolative view selection sampling the furthest target cameras. Our method achieves state-of-the-art performance in both extrapolative and interpolative settings. Qualitative results show our model effectively filters point cloud artifacts during warping, producing clean images with aligned geometry, while naive inpainting yields artifacts and inconsistent features.
@@
PixelSplat (Charatan et al., 2024) 14.01 0.582 0.384 23.85 0.806 0.185
MVSplat (Chen et al., 2024) 12.13 0.534 0.380 23.98 0.811 0.176
NoPoSplat (Ye et al., 2024) 14.36 0.538 0.389 25.03 0.838 0.160
@@
Table 2: In-domain comparison. We compare our model to existing feedforward NVS methods in Realestate10K (Zhou et al., 2018) dataset, our method being superior in extrapolative setting.
@@
In-domian results on RealEstate10K. Table 2 compares our method against feedforward approaches—PixelSplat (Charatan et al., 2024), MVSplat (Chen et al., 2024), DUSt3R (Wang et al., 2024), and NopoSplat (Ye et al., 2024)—on RealEstate10K (Zhou et al., 2018). We evaluate interpolative and extrapolative conditions using two views. For extrapolation, we sample references from the latter third and targets from the first third of video frames. Our model outperforms feedforward methods in extrapolative settings while maintaining competitive interpolative performance. Figure 6 shows conventional models struggle with large missing areas due to limited generative capabilities, whereas our approach effectively infers missing geometry and generates plausible imagery while preserving fidelity. We realistically inpaint partially visible objects with well-aligned geometry, attributed to geometric awareness from attention distillation regularizing structural attention maps.
Figure 7: Qualitative comparison with large model-based NVS methods. Qualitative comparison of our method to previous approaches demonstrates our method’s superior capabilities in conducting geometrically coherent image novel view synthesis with relatively short inference time.
In Fig 7, we additionally compare our method against recent large model-based novel view synthesis methods, namely LVSM (Jin et al., 2024), ZeroNVS (Sargent et al., 2023), and ViewCrafter (Yu et al., 2024). As shown in the figure, our method excels in extrapolative scenarios with significant unobserved regions. While LVSM accurately reconstructs areas with geometric overlap, it fails beyond observable boundaries, producing blurry content in occluded areas. ZeroNVS produces plausible novel views but lacks fidelity and requires manual specification of field-of-view, elevation, and content scale for each scene, with incorrect values leading to convergence failure. More critically, it requires approximately 2+ hours per scene for NeRF distillation via Score Distillation Sampling. ViewCrafter, evaluated with both 16-frame and 25-frame models, produces geometrically degraded imagery and artifacts under extrapolative settings, requiring 209.19 seconds for 25 frames on an A6000 GPU. In contrast, our generative approach leverages learned priors to synthesize plausible content for missing regions in only 9.67 seconds on average, maintaining high-fidelity detail and geometric consistency. This demonstrates our model’s competitive performance in both quality and efficiency in comparison to current large model-based novel view synthesis methods.
@@ 4.3 Ablation
(b) + Pointmap condition 16.93 0.594 0.243
(c) + Proximity-based mesh 17.01 0.601 0.238
(d) + Cross-modal instillation 17.41 0.614 0.229
@@
We conduct quantitative ablation studies on our method using the RealEstate10K dataset under the extrapolative setting. Our results in Table 3 demonstrate the contribution of each component. The baseline (a) receives no geometric conditioning, while (b) introduces naive pointmap conditioning. When we add proximity-based mesh conditioning (c) and cross-modal attention distillation (d), we observe progressive performance improvements. We provide additional qualitative results in Fig. 15 of our Appendix.
Analysis on the number of input viewpoints. As our model conducts aggregated attention to generate novel views from reference images, it can receive an arbitrary number of input viewpoints for generation. To demonstrate this, in Table 4 and Fig. 18 of our Appendix, we increase the number of reference viewpoints for a model trained at 2-viewpoint setting and analyze its effects in both image quality and geometric accuracy: the results demonstrate even without being trained on the given number of inputs, our model benefits strongly from additional viewpoints, showing the generalization capability of our aggregated attention architecture to various number of input reference viewpoints.
@@
4.4 Analysis on Robustness against External Geometry Predictors
[…]
No noise 15.580 0.615 0.184
Gaussian perturb. ($\sigma=3\%$) 14.778 0.520 0.213
Gaussian perturb. ($\sigma=6\%$) 14.501 0.507 0.225
Gaussian perturb. ($\sigma=10\%$) 14.129 0.487 0.239
Gaussian perturb. ($\sigma=15\%$) 13.726 0.465 0.262
Table 5: Robustness to Gaussian perturbation. Our model shows robustness against Gaussian perturbation to the correspondence condition.
[…]
No masking 15.580 0.615 0.184
30% masking 14.683 0.577 0.223
50% masking 13.610 0.468 0.272
80% masking 13.000 0.436 0.317
Table 6: Robustness to increase sparsity. Our model shows robustness against increased sparsity in the correspondence condition.
We demonstrate our model’s robustness to errors and artifacts from external geometry predictors by adding various perturbations to the correspondence conditions. First, we apply varying levels of Gaussian noise to predicted pointmap locations and show that our model exhibits minimal performance degradation even under high noise levels (Table 6). Second, we randomly mask points from the predicted pointmap to simulate sparse geometry, finding that our model maintains strong performance despite high sparsity in the correspondence condition (Table 6). Together, these results confirm that our diffusion-based framework’s generative capability effectively prevents error propagation from external priors, maintaining our pose-free methodology without compromising output quality. We provide qualitative evaluations of the same experiments in Fig. 10 and Fig. 11 of our Appendix.